Math 242 (Section 3)  University of South Carolina
Prof. Meade  Spring 1998

Supplemental Course Evaluation

Please return these forms in class on Thursday, April 23. You do not have to give your name, but if you wish to amplify your comments in person, please feel welcome to do so. Constructive criticism is appreciated! Thank you.

1. State three mathematical things that you have learned in the class so far.

(a)

(b)

(c)

2. State three non-mathematical things that you have learned in the class so far.

(a)

(b)

(c)

3. Do you feel the applied aspects of the course (word problems and projects) have helped you learn? Would you like to see more applications? If so, do you have any specific suggestions?

4. My impression is that when most of you use Maple you don’t initiate the commands yourself, but rather you re-use bits and pieces of the Maple code that I have distributed, making modifications as necessary. Is this impression correct? Do you think this is an effective way to use the software? Is there anything that could (or should) be done to facilitate your use of Maple (in or out of class)?

5. Concerning the integrated use of Maple in this class: did you find this aspect of the course enjoyable? did you find it educational? was this a more effective way of learning Maple than watching me give a demonstration? would you have preferred more or less use of the computer?